Back to the NIKA2 run 4 main page

March 2016, 15 (Tuesday)

March 2016, 14 (Monday)

March 2016, 13 (Sunday)

March 2016, 12 (Saturday)

March 2016, 11 (Friday)

March 2016, 10 (Thursday)

Early this morning we observed that some of the ethernet connections were not working and that one of the switch did not have power. After some investigation we asked for help to Juan Peñalver who concluded that there was no power on the device (direct measurement). Further we also realized after some tests that the cyrogenic amplifiers have no power neither. Unfortunately the telescope was in staw position and we could not reach the power supply unit. Finally, we also observe that some of the electronic boxes controlling the cryostat were not functioning. Thus, Juan P, and Juan-Luis took the decision to move the telescope in elevation for a few minutes.
After some investigation Juan P. found that a cable at the back (quite difficult access) of the box was not connected and reconnected it. This solved the problem. We restarted all the devices. The cryostat has warmed up. Andrea is right not having a careful look to it to restart the cool down. 

March 2016, 09 (Wednesday)

* We got the telescope at around 4am (Hervé and Alexandre)

* We took over in the morning around 9am (Juan and Anaëlle). Opacity around 0.35 at the begining, rising up to ~1 when we stopped at 3:30pm

March 2016, 08 (Tuesday)

We could not observe during the night because of bad weather. We started observations after maintenance. Unfortunately, there were first problems with PaKo and then with the telescop due to changes in mrt-lx1. After solving those problems we started normal observations but could not observe planets.


* We got the telescope at around 6pm after technical time and a few issues with the nika2 disks (disks full)

* We observe from 6pm to 10pm (Juan and Anaëlle) then will get the telescope back at 4am during the night (Hervé and Alexandre taking over)

* Opacity was good (about 0.2 @1mm at 6pm)

March 2016, 07 (Monday)


* Polarisation:

- 3X3 maps following the quasars 3C286, 3C279, 3C273 (X68)

- 1 skydip with HWP on the beam stopped

- 3X3 maps of NGC7538 (X12)

March 2016, 06 (Sunday)


* Intensity:

- 1 Beam maps (Uranus, telescope speed = 60arcsec/sec).

- 1 skydip without HWP on the beam

* Polarisation:

- 4X4 maps on Uranus (X4)

March 2016, 05 (Saturday)


* Intensity:

- 2 Beam maps (3C286, Mars).

- About 12 min integration on cluster RXJ1347-1145.

- 3X3 map on NGC7538 (X4 maps) with the HWP on the beam stopped

* Polarisation:

- 3X3 map on 3C286 (X4 maps)

- 3X3 map on NGC7538 (X6 maps)

- 16X16 maps on casA (splitted in three different maps)

- 4X4 maps on Uranus (X19)

- 2 skydips performed with HWP on the beam (the first with HWP on, the second with the HWP stopped)

March 2016, 04 (Friday)


- several maps on secondary calibrators (LKH234,NGC7027,CEPA,CRL2688,MWC349,K3-50A,G45.1,GL490,W3OH).

- Observation of NGC891 in intensity (@nkotf 20 20 -22.5 0 20 40 radec and @nkotf20 20 67.5 0 20 40 radec)

- 1 Beam maps on Uranus (telescope speeds - 78arcsec/s) and 2 beam maps on 3C279 (with telescope speeds - 70 and 60 arcsec/s).

- About 1h10 integration on cluster RXJ1347-1145.

- Second part of focus session.

We observed some secondary calibrators this night. Here are the maps of K3-50A and MWC349 we computed using a first analysis with nk_rta.



We also had the oportunity to observe RXJ1347 for 1h13min before the secondary calibrators became available. The following figure is the S/N map we get from the coaddition of the scans at 2 mm we reduced using a common_mode_kids_out decorrelation method (the dual_band_dec method leaves to much noise on the map so the S/N is reduced). The color bar is not diplayed because the noise properties have not been estimated yet. Therefore the S/N values are wrong. The SZ pic position is consistent with the one measured with the NIKA observations. Such a result is promising given the rather small integration time for an SZ observation !


We analyzed the beammaps made on Uranus yesterday morning at two different scan velocities (60 and 70"/sec). The beammaps were made at elevations higher than 30 degrees whereas this previous test realized on Mars on wednesday included a beammap measured at a mean elevation of 15 degree... The first three figures in the following show the results we get from the beammap scan made at 60"/sec (20160303s89_90_91) and the other three show the results we get from the beammap scan made at 70"/sec (20160303s86_87_88). It seems therefore than the scan velocity does not have a significant effect on the beam recovered shape and position in the focal plane. This conclusion has to be strengthened with more statistics so we will repeat this kind of observation tonight.

A1_Uranus_60.png A2_Uranus_60.png A3_Uranus_60.png

A1_Uranus_70.png A2_Uranus_70.png A3_Uranus_70.png

March 2016, 03 (Thursday)


We analyzed the first beammap we made yesterday on Mars (scan velocity of 60arcsec/s). As shown on the following figures, two feedlines on array 1 and 3 were very noisy so the recovered sensitivities are quite bad for these KIDs. We get 1014/1140 = 89% valid detectors on array 1, 847/1020 = 83% on array 2 and 918/1140 = 81% on array 3.

A1_sensitivity_Mars.png A2_sensitivity_Mars.png A3_sensitivity_Mars.png

We also analyzed the last beammap made on Uranus. As shown on the following figures there were no more noisy feedline but one electronic box on array 1 was not working during the acquisition. We therefore lost all the corresponding KIDs. We get 875/1140 = 77% valid detectors on array 1, 844/1020 = 83% on array 2 and 906/1140 = 79% on array 3.

A1_sensitivity_Uranus.png A2_sensitivity_Uranus.png A3_sensitivity_Uranus.png

March 2016, 02 (Wednesday)


* 02:00UT We go to the sky for the first time. Clear sky, opacity about 0.2 pretty stable.

The sky noise seems good on the 3 arrays (see the plots below). A little worse on array 3 (1mm NIKEL DC).




Figure. Noise on the sky for the 2mm, 1mm (array 1) and 1mm (array 3).

* The noise is 1.5Hz/sqrt(Hz), 2Hz/sqrt(Hz), and 3Hz/sqrt(Hz) for the 2mm, 1mm (array 1) and 1mm (array 3) respectively.

February 2016, 29 (Monday)

* Andrea arrived at the telescope.

* Status of the Instrument:

- The cryostat is at the nominam temperature T_100mK=150mK

- All the 20 feedlines are connected.

- The acquisition works nominally (few problems in the morning concerning the ports closed in nika2)

- We are producing normal,compressed and mini file on nika2 and mac12 during the night to check if they are right. First check is ok. We will keep checking tomorrow morning.

- We checked the modulation of the HWP and all the polarisation facilities of nika2. It seems to work nominally.

Hereafter some plot to check the good health of NIKA2:

run4_res_Array3.png run4_histo.png

Figure. Left: trace of the 1140 pixels of array 3; right: monitor with all the 20 boards tuned.

THE NOISE is taken with the window on the cryostat. It is in agreement with the previous observing campaign. No traces of the pulse tube.




Figure. Noise with the window on respectively the 2mm, 1mm NIKEL_RF (array 1) and 1mm NIKEL_DC (array 3)

DailyReportsNika2Run4 (last edited 2016-03-15 08:15:27 by NikaBolometer)